40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 735     Entry time: Thu Jul 24 19:29:26 2008
Author: Yoichi 
Type: Configuration 
Category: PSL 
Subject: C1:PSL-STAT_FSS_NOM_C_GAIN is changed from 30 to -0.7 
Koji, Yoichi

Since the light power going to the ref. cavity is now significantly increased (see Janne's elog later),
C1:PSL-STAT_FSS_NOM_C_GAIN
is changed from 30 to -0.7.
Otherwise, the MC did not lock.
ELOG V3.1.3-